# Fast Video Generation with Sliding Tile Attention

### v1

@@ Abstract
Diffusion Transformers (DiTs) with 3D full attention power state-of-the-art video generation, but suffer from prohibitive compute cost – when generating just a 5-second 720P video, attention alone takes 800 out of 945 seconds of total inference time. This paper introduces sliding tile attention (STA) to address this challenge. STA leverages the observation that attention scores in pretrained video diffusion models predominantly concentrate within localized 3D windows. By sliding and attending over the local spatial-temporal region, STA eliminates redundancy from full attention. Unlike traditional token-wise sliding window attention (SWA), STA operates tile-by-tile with a novel hardware-aware sliding window design, preserving expressiveness while being hardware-efficient. With careful kernel-level optimizations, STA offers the first efficient 2D/3D sliding-window-like attention implementation, achieving 58.79% MFU. Precisely, STA accelerates attention by 2.8–17× over FlashAttention-2 (FA2) and 1.6–10× over FlashAttention-3 (FA3). On the leading video DiT, HunyuanVideo, STA reduces end-to-end latency from 945s (FA3) to 685s without quality degradation, requiring no training. Enabling finetuning further lowers latency to 268s with only a 0.09% drop on VBench.
@@ 1 Introduction
Besides efficient computation, selecting optimal window sizes is crucial to preserving generation quality. We find that different attention heads exhibit specialized locality patterns – some heads focus on finer details in a small area, yet others capture broader context at a larger window – which we term as head specialization. Importantly, this head specialization remains agonistic to prompts, as evidenced in Figure 3. Based on this property, we develop a simple yet effective method to automatically configure the optimal window size per head via profiling, striking a balance between efficiency and quality. With STA, HunyuanVideo can generate a 5-second 720P video in 695s or 578s with no or minimal quality loss in a plug-and-play manner. In comparison, HunyuanVideo with FlashAttention-2 takes 1496s, while FlashAttention-3 takes 945s. STA achieves a end-to-end speedup of 2.15–2.59× over FlashAttention-2 and 1.35–1.63× over FlashAttention-3. Additionally, by fine-tuning diffusion models under more radical attention sparsity, we unlock even greater efficiency, delivering a 2.43 - 3.53$\times$ end-to-end speedup compared with FlashAttention-3.
@@ 3.2 Applying STA to Video Diffusion Model
0:  Transformer Model $M$, total steps $T$, Mask Pattern list $\mathcal{P}$, threshold $\delta$
0:  Dictionary $\mathrm{dict}$ that stores mask pattern result
@@
        for each $p$ in $\mathcal{P}$ (ordered by descending sparsity) do
@@
           $O^{\prime}\leftarrow$ (output of $M$ after masking)
           if $MSE(O,O^{\prime})<\delta$ then
              Record $m$ for $(h,l,t)$ in $\mathrm{dict}$
              break
@@
Training-free. As illustrated in Fig. 3, video diffusion models exhibit a pronounced 3D locality and head specialization pattern. Different transformer heads have different levels of locality, but their pattern is largely consistent across different prompts. We can exploit this property to search for the optimal window size for each head on a very small number of prompts, and expect the search result to work well on other prompts. We develop a simple heuristics to find such configuration in Algorithm 1. By adjusting the hyperparameter threshold $\delta$, we can balance the speedup and the quality loss.
@@ 4.2 Human Evaluations
Figure 7: Human evaluation on 200 prompts from the MovieGen Bench (Polyak et al., 2024). STA achieves a 1.36× end-to-end speedup while maintaining performance comparable to the original HunyuanVideo. Additionally, STA consistently outperforms $\Delta$-DiT across different inference budgets.
We assess human preference across five models that achieve the best quality performance: (1) HunyuanVideo; (2) STA-tf-1.36x: HunyuanVideo with 1.36× speedup via training-free mask search, (3) STA-t-2.43x: HunyuanVideo with 2.43× speedup via finetuning with STA, (4-5) two variants of $\Delta$-DiT (1.36× and 1.8× speedup). Other baselines such as CLEAR or Swin are either prohibitively slow or produce subpar quality. Following MovieGen (Polyak et al., 2024), we randomly sample 200 prompts from the MovieGen Bench and conduct pairwise comparisons between these models. Evaluators select the video with higher overall quality or mark both as a tie.
In Figure 7, STA-t-2.43x decisively outperforms $\Delta$-DiT-1.8x, achieving a dominant 69.8% win rate versus 11.1%, despite a greater speedup. Similarly, STA-tf-1.36x surpasses $\Delta$-DiT-1.36x with a 58.5% win rate against 11.0%. Compared to the original HunyuanVideo, STA maintains competitive quality, with STA-tf-1.36x achieving a 74.5% tie rate—indicating near-parity in most cases. Though it has a 5.5 percentage point lower win rate than its loss rate, this tradeoff comes with a 1.36× speedup, demonstrating strong quality preservation alongside significant efficiency gains. These results establish STA as achieving a superior quality-efficiency tradeoff compared to $\Delta$-DiT.
@@ 4.3 Training-free Results
In Table 3, we evaluate mask-search STA  and $\Delta$-DiT on VBench prompts, testing robustness across different sampling steps. Our focus is on preserving video quality in a low-step regime. For each diffusion step count, we report SSIM, PSNR, and CD-FVD, using HunyuanVideo’s outputs at the same step count as the reference. We set the scheduler shift to 17 for 10-step inference and 7 for 25- and 50-step inference following Hunyuan’s default settings.
Table 3: Training-free performance with varying sampling steps. $\Delta$-DiT shows consistently worse quality compared to STA, and that gap widens as the number of inference steps decrease.
@@
$\Delta$-DiT 72.86% 18.09 122.74 693s 1.36×1.36$\times$1.36 ×
STA 76.21% 19.94 97.03 695s 1.36×1.36$\times$1.36 ×
@@
$\Delta$-DiT 77.91% 19.86 196.25 352s 1.34×1.34$\times$1.34 ×
STA 82.47% 22.53 95.86 348s 1.36×1.36$\times$1.36 ×
@@
$\Delta$-DiT 83.19% 21.20 201.24 144s 1.32×1.32$\times$1.32 ×
STA 87.15% 24.04 80.41 139s 1.36×1.36$\times$1.36 ×
Our training-free STA consistently outperforms $\Delta$-DiT, with the quality gap widening as the step count decreases. At 50 steps, $\Delta$-DiT’s CD-FVD score is 25.71 higher than STA (122.74 vs. 97.03, where lower is better). This gap grows to 100.39 at 25 steps and 120.83 at 10 steps. Qualitatively, $\Delta$-DiT produces structurally similar but visually degraded outputs, explaining the narrow SSIM gap—SSIM captures structural similarity but not fine details. These results suggest that caching-based methods degrade in very low-step sampling, while STA maintains fidelity to the original model. Intuitively, caching methods exploit temporal redundancy, which diminish with fewer steps, whereas STA remains effective. This distinction suggests STA may complement step-reduction techniques like consistency distillation (Song et al., 2023), which we leave for future work.
@@ 5 Related Work
We review additional related work in diffusion acceleration. Linear attention methods (Wang et al., 2020; Liu et al., 2021a; Arar et al., 2022; Yang et al., 2024a) can decompose the softmax operation in quadratic attention using kernel or gate functions to achieve linear complexity. However, these methods have not yet been successful in video DiTs. Another major bottleneck in diffusion models is the large number of diffusion steps. Several techniques have been proposed to accelerate sampling without sacrificing quality, including DDIM (Song et al., 2020) and faster ODE and PDE solvers using numerical methods (Song & Ermon, 2019; Lu et al., 2022a, b; Jolicoeur-Martineau et al., 2021). New methods have also emerged to further reduce the number of sampling steps, such as consistency distillation (Kim et al., 2023; Song et al., 2023; Salimans et al., 2024; Xie et al., 2024), adversarial distillation (Sauer et al., 2023), and other distillation approaches (Li et al., 2024; Yin et al., 2023, 2024). STA is largely complementary to these methods.

### v2

@@ Abstract
Diffusion Transformers (DiTs) with 3D full attention power state-of-the-art video generation, but suffer from prohibitive compute cost – when generating just a 5-second 720P video, attention alone takes 800 out of 945 seconds of total inference time. This paper introduces sliding tile attention (STA) to address this challenge. STA leverages the observation that attention scores in pretrained video diffusion models predominantly concentrate within localized 3D windows. By sliding and attending over the local spatial-temporal region, STA eliminates redundancy from full attention. Unlike traditional token-wise sliding window attention (SWA), STA operates tile-by-tile with a novel hardware-aware sliding window design, preserving expressiveness while being hardware-efficient. With careful kernel-level optimizations, STA offers the first efficient 2D/3D sliding-window-like attention implementation, achieving 58.79% MFU. Precisely, STA accelerates attention by 2.8–17× over FlashAttention-2 (FA2) and 1.6–10× over FlashAttention-3 (FA3). On the leading video DiT, HunyuanVideo, STA reduces end-to-end latency from 945s (FA3) to 501s without quality degradation, requiring no training. Enabling finetuning further lowers latency to 268s with only a 0.09% drop on VBench. We make our codebase public at https://github.com/hao-ai-lab/FastVideo.
@@ 1 Introduction
Besides efficient computation, selecting optimal window sizes is crucial to preserving generation quality. We find that different attention heads exhibit specialized locality patterns – some heads focus on finer details in a small area, yet others capture broader context at a larger window – which we term as head specialization. Importantly, this head specialization remains agonistic to prompts, as evidenced in Figure 3. Based on this property, we develop a simple yet effective method to automatically configure the optimal window size per head via profiling, striking a balance between efficiency and quality. With STA, HunyuanVideo can generate a 5-second 720P video in 501s with no or minimal quality loss in a plug-and-play manner. In comparison, HunyuanVideo with FlashAttention-2 takes 1496s, while FlashAttention-3 takes 945s. STA achieves a end-to-end speedup of 2.98× over FlashAttention-2 and 1.89× over FlashAttention-3. Additionally, by fine-tuning diffusion models under more radical attention sparsity, we unlock even greater efficiency, delivering a 2.43 - 3.53$\times$ end-to-end speedup compared with FlashAttention-3.
@@ 3.2 Applying STA to Video Diffusion Model
0:  Transformer model $M$, Total steps $T$, Mask pattern list $\mathcal{P}$, Keep first $T_{0}$ timestep full attn
0:  Dictionary $\mathrm{dict}$ that stores selected mask pattern for each head
@@
        $O\leftarrow$ (attn output of original $(l,h)$ )
        Initialize $\mathrm{minimum\_loss}\leftarrow\infty$
        Initialize $\mathrm{best\_pattern}\leftarrow$ null
        for each $p$ in $\mathcal{P}$ do
@@
           $O^{\prime}\leftarrow$ (attn output of $M$ after masking)
           $\mathrm{loss}\leftarrow MSE(O,O^{\prime})$
           if $\mathrm{loss}<\mathrm{minimum\_loss}$ then
              $\mathrm{minimum\_loss}\leftarrow\mathrm{loss}$
              $\mathrm{best\_pattern}\leftarrow p$
        Record $\mathrm{best\_pattern}$ for $(t,l,h)$ in $\mathrm{dict}$
@@
Training-free. As illustrated in Fig. 3, video diffusion models exhibit a pronounced 3D locality and head specialization pattern. Different transformer heads have different levels of locality, but their pattern is largely consistent across different prompts. We can exploit this property to search for the optimal window size for each head on a very small number of prompts, and expect the search result to work well on other prompts. We develop a simple heuristics to find such configuration in Algorithm 1 and decide the final configuration by averaging the mask-search loss across 16 prompts. In practice, we keep full attention for the initial $T_{0}$ timesteps  (Li et al., 2024a; Zhao et al., 2024; Lv et al., 2024), and then apply STA for the rest timesteps.
@@ 4.2 Human Evaluations
Figure 7: Human evaluation on 200 prompts from the MovieGen Bench (Polyak et al., 2024). STA achieves a 1.89× end-to-end speedup while maintaining performance comparable to the original HunyuanVideo. Additionally, STA consistently outperforms $\Delta$-DiT across different inference budgets.
We assess human preference across five models that achieve the best quality performance: (1) HunyuanVideo; (2) STA-tf-1.89x: HunyuanVideo with 1.89× speedup via training-free mask search, (3) STA-t-2.43x: HunyuanVideo with 2.43× speedup via finetuning with STA, (4-5) two variants of $\Delta$-DiT (1.36× and 1.8× speedup). Other baselines such as CLEAR or Swin are either prohibitively slow or produce subpar quality. Following MovieGen (Polyak et al., 2024), we randomly sample 200 prompts from the MovieGen Bench and conduct pairwise comparisons between these models. Evaluators select the video with higher overall quality or mark both as a tie.
In Figure 7, STA-t-2.43x decisively outperforms $\Delta$-DiT-1.8x, achieving a dominant 70.0% win rate versus 11.0%, despite a greater speedup. Similarly, STA-tf-1.89x surpasses $\Delta$-DiT-1.36x with a 66.5% win rate against 10.0%. Compared to the original HunyuanVideo, STA maintains competitive quality, with STA-tf-1.89x achieving a 83.0% tie rate—indicating near-parity in most cases. Though it has a 7.0 percentage point lower win rate than its loss rate, this tradeoff comes with a 1.89× speedup, demonstrating strong quality preservation alongside significant efficiency gains. These results establish STA as achieving a superior quality-efficiency tradeoff compared to $\Delta$-DiT.
@@ 4.3 Training-free Results
In Table 3, we evaluate mask-search STA  and $\Delta$-DiT on VBench prompts, testing robustness across different sampling steps. For each diffusion step count, we report SSIM, PSNR, and CD-FVD, using HunyuanVideo’s outputs at the same step count as the reference. We set the scheduler shift to 17 for 10-step inference and 7 for 25- and 50-step inference following Hunyuan’s default settings. We keep full attn for the first 12,6,3 steps for 50-step, 25-step, and 10-step inference respectively.
Table 3: Training-free performance with varying sampling steps. $\Delta$-DiT shows consistently worse quality compared to STA
@@
$\Delta$-DiT 72.86 18.09 122.74 693s 1.36×1.36$\times$1.36 ×
STA 87.67 28.76 66.12 501s 1.89×1.89$\times$1.89 ×
@@
$\Delta$-DiT 77.91 19.86 196.25 352s 1.34×1.34$\times$1.34 ×
STA 88.96 28.99 76.34 250s 1.89×1.89$\times$1.89 ×
@@
$\Delta$-DiT 83.19 21.20 201.24 144s 1.32×1.32$\times$1.32 ×
STA 87.84 27.14 84.80 105s 1.76×1.76$\times$1.76 ×
Our training-free STA consistently outperforms $\Delta$-DiT even with much higher speedup. At 50 steps, STA achieves substantial improvements in all metrics, with 14.81 higher SSIM (87.67 vs. 72.86) and 10.67 higher PSNR (28.76 vs. 18.09). $\Delta$-DiT’s CD-FVD score is 56.62 higher than STA (122.74 vs. 66.12, where lower is better). This gap grows to 119.91 at 25 steps and 116.44 at 10 steps. Qualitatively, $\Delta$-DiT consistently produces visually degraded outputs across all sampling steps, exhibiting compromised structural similarity and diminished fine details, while STA maintains high fidelity to the original model.
@@ 5 Related Work
We review additional related work in diffusion acceleration. Linear attention methods (Wang et al., 2020; Liu et al., 2021a; Arar et al., 2022; Yang et al., 2024a) can decompose the softmax operation in quadratic attention using kernel or gate functions to achieve linear complexity. However, these methods have not yet been successful in video DiTs. Another major bottleneck in diffusion models is the large number of diffusion steps. Several techniques have been proposed to accelerate sampling without sacrificing quality, including DDIM (Song et al., 2020) and faster ODE and PDE solvers using numerical methods (Song & Ermon, 2019; Lu et al., 2022a, b; Jolicoeur-Martineau et al., 2021). New methods have also emerged to further reduce the number of sampling steps, such as consistency distillation (Kim et al., 2023; Song et al., 2023; Salimans et al., 2024; Xie et al., 2024), adversarial distillation (Sauer et al., 2023), and other distillation approaches (Li et al., 2024b; Yin et al., 2023, 2024). STA is largely complementary to these methods.